Sketch the graph of f: x->x+|x+2| where x€R


1
Expert's answer
2022-04-22T02:32:43-0400

"f(x)=x+|x+2|=\begin{cases}
x+x+2,\ \ \text{if} \ x\geq -2
\\
x-(x+2),\ \ \text{if}\ x<-2
\end{cases}
=\begin{cases}
2x+2,\quad \text{if}\ \ x\geq -2
\\
-2,\qquad \ \ \text{if}\ \ x<-2
\end{cases}"